Lilian9667:

-4+x=6 A.) -24 B.) -10 C.)-2 D.) 10

Lilian9667:

@MiraAngel

Lilian9667:

I need help with this

MiraAngel:

add 4 on both sides to find x -4+x=6 +4 +4 x=10 Answer is D I hope this helps! Let me know if you need anymore help! :) Good Luck!

Lilian9667:

Thank you! It makes sense now

MiraAngel:

Great :)